OpenStudy (anonymous):

Can someone talk me through this problem and explain the process? A solution is made by dissolving 1.00 moles of sodium chloride (NaCl) in 155 grams of water. If the molal boiling point constant for water (Kb) is 0.51 °C/m, what would be the boiling point of this solution?

OpenStudy (aaronq):

1. Find the molality \(m=\dfrac{n_{solute}}{kg_{solvent}}\) 2. substitute the values given into: \(\Delta T =i*m*K_b\) i = van't hoff factor; approximated by the number of molecules/atoms present upon dissociation of species in the solvent. NaCl (s) -> \(Na^+_{\;aq.} +Cl^-_{\;aq.}\) i=2 \(\Delta T\) = change in temperature m=molality Kb= boiling point elevation constant After you found \(\Delta T\), you add it to the temp at which the solvent boils. Water boils at 100 Celsius, so, for this question: Boiling Temp of solution = \(100 C^{\;o} + \Delta T\)

OpenStudy (anonymous):

Thank you! This helped me so much!

OpenStudy (aaronq):

no problem, glad i could help !